OpenStudy (anonymous):

Write an equation for the circle that satisfies the given set of conditions. 1. Center (8, -3), radius 6 2. Center (5, -6), radius 4 3. Center (-5, 2), Passes through (-9, 6) 4. Endpoints of a diameter at (6, 6) and (10, 12)

OpenStudy (anonymous):

1.(x-8)^2+(y+3)^2=6^2

OpenStudy (anonymous):

(x1,y1)is the centre and r is the radius

OpenStudy (anonymous):

(x-8)^2 +(y+3)^2 =36

OpenStudy (anonymous):

(x-8)^2 +(y+3)^2 =36

OpenStudy (anonymous):

That didnt help me what so ever

ganeshie8 (ganeshie8):

hey lets do from scratch

OpenStudy (anonymous):

Thank you!

ganeshie8 (ganeshie8):

1. Center (8, -3), radius 6

OpenStudy (anonymous):

alright

ganeshie8 (ganeshie8):

we use this :- equation of circle wid center \((h, k)\) and radius \(r\) :- \((x-h)^2 + (y-k)^2 = r^2\)

ganeshie8 (ganeshie8):

equation of circle wid center (8, -3) and radius 6 :- \((x-8)^2 + (y--3)^2 = 6^2\) \((x-8)^2 + (y+3)^2 = 36\)

OpenStudy (anonymous):

okay. That makes sense

ganeshie8 (ganeshie8):

cool :)

ganeshie8 (ganeshie8):

try next q for urself :) il correct if u go wrong, give it a try

ganeshie8 (ganeshie8):

can u ? :)

OpenStudy (anonymous):

\[(x-5)^{2}+(y+6)^{2}=4^{2}\] \[(x-5)^{2}+(y+6)^{2}=16\]

OpenStudy (anonymous):

sorry my computer was bugging out for a second

OpenStudy (anonymous):

What equation do i use for the other two problems

ganeshie8 (ganeshie8):

ping once u r back :)

OpenStudy (anonymous):

i am